OpenStudy (anonymous):

1 3/4 x 23 6/5 + 16 =

OpenStudy (turingtest):

unmix your mixed numbers: (7/4)(121/5)+16=847/20+320/20=1167/20=

OpenStudy (turingtest):

1167/20 that's all

OpenStudy (anonymous):

okkkk thxxxs